Detailed Explanation of the following answer is required if we have given the problem as
If the total revenue function is TR=100,000Q-10Q^2 and the total cost function is TC=10,000+20Q+Q^2, determine
A) the output rate that will maximize total royalty revenue and the amount of royalty income hat smith and Wesson would receive. B) the output rate that would maximize profit to the publisher.
Based on this rate of output, what is the amount of royalty income smith and Wesson would receive? Compare the royalty income of smith and Wesson to that determined in part
Solution is given as below:
A) TR will be maximized if Q = 0, then TR = 100,000 and royalty = 100,000 x 0,15 =

=15,000

B) TC will be minimum if Q = 0, then TC = 90,000 and royalty = 100,000 x 0,15 =

= 15,000.
1
Expert's answer
2015-07-22T00:00:44-0400
If the total revenue function is TR = 100,000Q - 10Q^2 and the total cost function is TC = 10,000 + 20Q + Q^2, determine
A) the output rate that will maximize total royalty revenue and the amount of royalty income hat smith and Wesson would receive is:
TR will be maximized if MR = 0, then MR = TR' = 100,000 - 20Q
TR = 100,000 and royalty = (100,000 - 20Q) x 0,15
B) the output rate that would maximize profit to the publisher.
TC will be minimum if Q = 0, then TC­ = 90,000 and royalty = 100,000 x 0,15 = 15,000.